GENERALIZED CALIBRATION APPROACH FOR ESTIMATING VARIANCE IN SURVEY SAMPLING

SARJINDER SINGH

Department of Mathematics and Statistics, University of Windsor, Windsor, Ontario, Canada N9B 3P4

(Received January 27, 1999; revised July 21, 1999)

Abstract. In the present investigation, a general set-up for inference from survey data that covers the estimation of variance of estimators of totals and distribution functions has been considered, using known higher order moments of auxiliary information at the estimation stage. Several estimators of variance of estimators of totals and distribution functions are shown to be the special cases of the proposed strategy. An empirical study has also been given to show the performance of the proposed estimators over the existing estimators in the literature.

Key words and phrases: Auxiliary information, distribution functions, totals, estimation of variance.